OpenStudy (anonymous):

A polygon has an area of 144 square inches and one of its sides is 8 inches long. If a second similar polygon has an area of 64 square inches, what is the length of the corresponding side in the second polygon?

Directrix (directrix):

If two polygons are similar, then the square of the scale factor of the two polygons is the ratio of the areas

Directrix (directrix):

Let x be the side of the second polygon which corresponds to the side of length 8 in the first polygon.

Directrix (directrix):

(8/x)^2 = 144/64

Directrix (directrix):

8/x = 12/8

OpenStudy (anonymous):

oh ok, thanks, that make much more since thanks again!

Directrix (directrix):

Glad to help. Check my arithmetic.